# Clear Minds Think Alike: What Makes LLM Fine-tuning Robust? A Study of Token Perplexity

### v1

@@ Abstract
Maintaining consistent model performance across domains is a fundamental challenge in machine learning. While recent work has explored using LLM-generated data for fine-tuning, its impact on cross-domain generalization remains poorly understood. In this paper, we present a systematic analysis revealing that fine-tuning with LLM-generated data not only improves target task performance but also reduces out-of-domain (OOD) degradation compared to fine-tuning with ground truth data. Through analyzing the data sequence in tasks of various domains, we demonstrate that this enhanced OOD robustness stems from a reduced prevalence of high perplexity tokens in LLM-generated sequences. Following this hypothesis we showed that masking high perplexity tokens in ground truth training data also achieves similar OOD preservation comparable to using LLM-generated data. Extensive experiments across diverse model architectures and scales, including Gemma2-2B, Mistral-7B and Llama3-8B, corroborate the consistency of our findings. To the best of our knowledge, this work provides the first mechanistic explanation for the superior OOD robustness conferred by LLM-generated training data, offering valuable insights for developing more robust fine-tuning strategies.
@@ 1 Introduction
Supervised fine-tuning (SFT) on Large Language Models (LLMs), particularly for instruction following [20], have demonstrated remarkable capabilities in comprehending novel instructions [31] and generating useful outputs across various tasks, including summarization and web queries [26, 18, 1, 12]. As LLM applications continue to expand into specialized domains such as data querying and programming assistance [28, 19], practitioners face significant computational and data constraints when adapting LLMs. This challenge is further compounded by the performance saturation observed in smaller LLMs, where benchmark scores on tasks like MATH [10] and MBPP[2] plateau under limited parameter sizes, probably due to inherent limitations in model capacity. Fine-tuning such models often risks degrading their general capabilities [23, 16, 22]. A critical challenge emerges: how to efficiently fine-tune pre-trained instruction-following LLMs for domain-specific tasks while preserving their general capabilities, especially when the original fine-tuning dataset is inaccessible or when computational resources are limited.
@@
Figure 1: An example MATH problem showing high perplexity tokens (highlighted in red, perplexity $\geq 2.5$) in both ground truth and Self-Output responses.
Recent work has demonstrated the effectiveness of using model-generated data for fine-tuning, showing that training on high-quality samples from model generations [36, 8, 24] can outperform training on ground truth data. While Ren et al. [23] showed that model-generated responses yield superior performance compared to human annotations, their analysis focused on using more capable models (e.g. GPT-4) to generate training data and relieve the degradation of general tasks. However, although they mange to use "familiarity", or low perplexity, of generated response to explain how unfamiliar data causes lower performance, it is still not fully understood that how generated data works well for instruction tuning: (1) Context of regenerated response differs from ground truth response so the bias in context could disturb the effectiveness of low perplexity training in their experiments. (2) the impact of low perplexity training on out-of-domain performance is still poorly understood, and (3) The distillation of larger-sized model generally benefits the training, it is better to eliminate the effect of larger-sized model and discuss the low perplexity training issue with the same models. Our work provides the first mechanistic explanation for why LLM-generated data leads to more robust fine-tuning outcomes. Through systematic analysis, first, we demonstrate that fine-tuning with LLM-generated data significantly reduces out-of-domain degradation compared to ground truth data training. Second, we trace this improvement to a key characteristic as Figure 1 shows. The LLM generated response has lower sentence perplexity and lower proportion of high perplexity tokens. Hence, this finding leads to a practical insight: simply masking high perplexity tokens in ground truth training data achieves comparable out-of-domain preservation to using LLM-generated data for training. In such way, we also manage to use the same model before training to eliminate high perplexity tokens from the original ground truth context, and discover the improvement that benefits from low perplexity training.
Our contributions are threefold:
  • We identify the mechanistic explanation for out-of-domain performance degradation through our proposed method: Selective Token Masking (STM), a simple yet effective training strategy to filter out high perplexity tokens to achieve low perplexity training.
  • We provide the first empirical evidence that LLM-generated training data significantly reduces out-of-domain degradation during fine-tuning.
  • We demonstrate the broad applicability of these findings through comprehensive experiments across different model architectures and scales, including Gemma2-2B, Mistral-7B, and Llama3-8B.
These insights offer new directions for developing robust fine-tuning strategies that better preserve general capabilities while enhancing target task performance.
@@ 2 Training LLMs on Self-Generated Data
To examine the effectiveness of LLM-generated data for fine-tuning, we investigate two distinct methods for generating training data of two different tasks and evaluate the models trained with the generated data on five different tasks. Together, these two methods, Self-Output and Rephrase, provide a framework for generating LLM-based training data, each addressing different challenges and trade-offs. In this section, we outline our original datasets for LLMs to generate from, and the approaches to creating self-generated training datasets with language models.
@@ 2.1 Training and Evaluation Framework
For training in-domain task, we adopt MBPP and MATH datasets for training. These two datasets provide complete guidance and comprehensive ground truth labels for models to learn from, including complete solutions, reasoning paths, and test cases, not just final answers. For evaluation, besides the testing set of MBPP and MATH, we also evaluate on GSM8k [5], ARC-Challenge [4] and BIRD [14] to evaluate fine-tuned models performance in different aspects of generation. Detailed of dataset will be introduced in Section 4.
@@ 2.2 Self-Output
Self-Output method follows the same high level concept as other synthetic generation works such as [36, 23] where we start by sampling a diverse LLM response via strict filtering process thereby creating a high-quality synthetic dataset. For each training instance, we use a language model $M$ such as Llama3-8B Instruct to generate $N$ (e.g. 32) distinct responses with temperature $T=0.7$. We then filter the generated responses which are semantically aligned with the ground truth to ensure the quality of our synthetic dataset.
A key advantage of Self-Output is its ability to produce reliable training data when the ground truth is verifiable, such as in programming tasks or other domains with unambiguous correct answers. However, this method is limited to scenarios where such verification is feasible, as it relies on the availability of a clear ground truth for filtering.
Table 2 presents the performance comparison between Self-Output, Rephrase, Ground Truth training and base models across both in-domain and out-of-domain tasks using different models in terms of model size and model series. The base models in this paper refer to generic instruction following models before training, e.g. Llama3-8B-Instruct [6], Gemma2-2B Instruct [27] and Mistral-7B Instruct v0.3 [11]. For in-domain performance, Self-Output demonstrates consistent improvements in both MBPP and MATH compared to their respective Ground Truth baselines. For out-of-domain scenarios, the Self-Output method achieves robust performance compared with the base models in most cases, and some could improve. We also observe a particularly strong transfer from MATH training to GSM8K evaluation, though performance on the BIRD dataset remains challenging across all methods. These results suggest that our Self-Output method enhances model generalization, particularly in mathematical reasoning tasks, while maintaining competitive performance in programming and general knowledge domains. For more details, such as performance comparison with other different models, please refer to Appendix A.
@@ 2.3 Rephrase
Yang et al. [35] proposed self-distillation. The process uses an instruction-finetuned LLM to rephrase ground-truth responses in its own style. By providing both the instruction and ground truth to the target LLM, it generates semantically equivalent reformulations with ground truth response. Compared to Self-Output, Rephrase is computationally efficient, as it requires only a single pass of generation without the need for additional filtering. However, Rephrase is susceptible to hallucinations, as evidenced by the higher variance in token level perplexity (see Table 1). Despite this limitation, Rephrase offers versatility, as it does not rely on verifiable ground truth and can be applied to a broader range of tasks. We apply Rephrase as one of the baselines in Table 2.
Together, these methods provide a framework for generating LLM-based training data, each addressing different challenges and trade-offs. In the following section, we analyze the perplexity of datasets generated by these methods to quantify their impact on model uncertainty and fine-tuning stability. These insights offer new directions for developing robust fine-tuning strategies that better preserve general capabilities while enhancing target task performance.
2.4 Analysis of Perplexity Across Each Samples
In this section, we analyze the perplexity of datasets generated by Self-Output and Rephrase to quantify their impact on model uncertainty and fine-tuning stability. These insights offer new directions for developing robust fine-tuning strategies that better preserve general capabilities while enhancing target task performance. We first collect the Llama3-8B Instruct inference statistics from all three methods: Ground Truth, Rephrase and Self-Output strategy. We evaluate the perplexity of responses by summing up the negative log probability of each token conditioned on its previous context in the sentence to measure the uncertainty of model outputs over the entire sequence. Specifically, the perplexity $PPL$ [^1] In this section we use natural log instead of base 2 for computation convenient across a sequence is computed as:
$PPL=\exp(\frac{1}{N}\sum_{i=1}^{N}P(w_{i}|w_{1},...,w_{i-1}))$ (1)
where $P(w_{i}|w_{1},...,w_{i-1})$ represents the conditional probability of token $w_{i}$ given its preceding context, and $N$ is the total number of tokens in the sequence.
Table 1: Perplexity average and variance score over answer sequences of MBPP and MATH training datasets on Llama3-8B Instruct. The average sentence perplexity are calculated by averaging the sum of each sentence’s perplexity over token perplexity score.
Data Method Average Perplexity
MBPP Ground Truth 4.83 (7.04)
Rephrase 1.69 (0.16)
Self-Output 1.16 (0.01)
MATH Ground Truth 2.45 (0.81)
Rephrase 2.11 (9.28)
Self-Output 1.34 (0.03)
Table 1 demonstrates that the Self-Output strategy achieves both the lowest average perplexity of 1.16. Among the three approaches, Ground Truth exhibits the highest uncertainty with mean perplexity of 4.83, and the perplexity of Rephrase are slightly higher than Self-Output. The consistently lower uncertainty metrics of Self-Output suggest that LLMs are less likely to deviate far from their initial weights when fine-tuning on downstream tasks created by Self-Output. This characteristic maintains model original performance on out-of-domain tasks while effectively learning from the in-domain task training data.
@@ 2.4 Analysis of Perplexity Across Each Samples
Figure 2: Comparison of token-level perplexity (PPL) distributions between human-annotated ground truth (left) and Llama3-8B generations for Rephrase and Self-Output sequence (middle and right).
@@
To investigate token-level predictability we sample a single MATH response and visualize the token level perplexity value. Figure 2 reveals distinct patterns across Ground Truth, Rephrase, and Self-Output responses. The first 20 tokens variance are 2.04 for Ground Truth, 2.31 for Rephrase, and 0.04 for Self-Output which confirm our observation: Ground Truth exhibits frequent perplexity spikes of high-amplitude (perplexity around 6-8), which reflects the natural variability in human problem-solving expressions. While Rephrase maintains similar peak amplitudes to Ground Truth, i.e. more regular spike patterns, it suggests a hybrid variability between human and LLMs. In contrast, Self-Output demonstrates remarkably consistent low token perplexity values rarely exceeding 2, indicating highly predictable token sequences. These patterns suggest that Self-Output generates more deterministic solutions, potentially beneficial for maintaining consistent problem-solving approaches, while human responses introduce natural variations.
To demonstrate the different content of tokens distribution in between Ground Truth and Self-output responses, please refer to Appendix B.
3 Selective Token Masking (STM)
@@
Previous analysis of Figure 2 reveals a significant discrepancy in token perplexity distribution between Self-Output generated data and other approaches. This observation raises a fundamental question: Is the superior performance of Self-Output methods directly attributable to the lower perplexity of their generated tokens? To investigate this hypothesis and develop more efficient fine-tuning strategies, we propose Selective Token Masking (STM), a novel approach for supervised fine-tuning (SFT) that focuses on the role of token perplexity in model training.
3.1 Selective Token Masking
@@
The core mechanism of STM is straightforward yet effective: using an existing instruction-tuned model to compute token perplexities, we mask and exclude ground truth tokens that exceed a predefined threshold $\tau$ during the fine-tuning process. We investigate the foundations of successful fine-tuning through two key findings. First, STM reveals that token perplexity plays a crucial determinant of fine-tuning outcomes. Second, STM demonstrates that the performance gains of Self-Output methods can be achieved through direct perplexity-based token management, suggesting these gains stem from low perplexity token distributions rather than the self-generation process itself. This approach not only provides computational efficiency but also establishes a systematic framework for understanding model adaptation mechanisms.
3.2 Comparison
@@
Our approach shares some conceptual similarities with [15], yet differs significantly in implementation and efficiency. Their method requires a two-stage process: first training a reference model on high quality data to identify high perplexity tokens from ground truth data as learnable tokens, and then pre-trains an LLM on the learnable tokens. In contrast, STM achieves similar objectives through a streamlined single-stage process: it directly utilizes an existing instruction-tuned model to compute token perplexities via a single forward pass. Thus, STM eliminate the computational overhead of [15], while sustaining to the performance of low perplexity training.
4 Experiment Setting
To evaluate the effectiveness of our proposed method, we select benchmark datasets with train sets designed to evaluate the robustness on the in-domain and out-of-domain data after fine-tuning. We only chose datasets with a verifiable ground truth such that we can filter incorrect self-generated responses out for better data quality.
Table 2: Performance for Ground Truth, Rephrase, Self-Output, and STM methods trained on datasets MBPP and MATH respectively. Values of a cell block in gray represent the in-domain testing performance of the model trained with in-domain task data.
Model in-domain method MBPP MATH ARC gsm8k BIRD
task
Gemma2-2B - base model 51.06 28.10 74.74 58.45 13.89
MBPP ground truth 39.95 $\downarrow$ 22.90 $\downarrow$ 71.59 $\downarrow$ 19.26 $\downarrow$ 5.14 $\downarrow$
Rephrase 48.41 $\downarrow$ 28.90 $\uparrow$ 70.56 $\downarrow$ 48.22 $\downarrow$ 13.75 $\downarrow$
Self-Output 53.44 $\uparrow$ 29.20 $\uparrow$ 58.45 $\downarrow$ 52.39 $\downarrow$ 13.95 $\uparrow$
$STM_{\tau=2.5}$ 49.21 $\downarrow$ 29.80 $\uparrow$ 74.83 $\uparrow$ 56.41 $\downarrow$ 13.75 $\downarrow$
MATH ground truth 45.77 $\downarrow$ 21.70 $\downarrow$ 29.52 $\downarrow$ 19.03 $\downarrow$ 12.45 $\downarrow$
Rephrase 46.83 $\downarrow$ 20.00 $\downarrow$ 73.38 $\downarrow$ 37.00 $\downarrow$ 12.58 $\downarrow$
Self-Output 50.00 $\downarrow$ 30.90 $\uparrow$ 72.95 $\downarrow$ 57.01 $\downarrow$ 13.89 -
$STM_{\tau=2.5}$ 49.47 $\downarrow$ 30.30 $\uparrow$ 75.34 $\uparrow$ 55.42 $\downarrow$ 13.30 $\downarrow$
Llama3-8B - base model 59.26 31.40 78.16 72.55 20.53
MBPP ground truth 57.94 $\downarrow$ 23.30 $\downarrow$ 59.13 $\downarrow$ 18.50 $\downarrow$ 17.60 $\downarrow$
Rephrase 54.06 $\downarrow$ 30.00 $\downarrow$ 75.43 $\downarrow$ 72.40 $\downarrow$ 19.16 $\downarrow$
Self-Output 61.64 $\uparrow$ 33.40 $\uparrow$ 79.18 $\uparrow$ 78.39 $\uparrow$ 20.01 $\downarrow$
$STM_{\tau=2.5}$ 61.11 $\uparrow$ 32.00 $\uparrow$ 78.92 $\uparrow$ 75.43 $\uparrow$ 20.66 $\uparrow$
MATH ground truth 58.73 $\downarrow$ 25.80 $\downarrow$ 48.63 $\downarrow$ 67.93 $\downarrow$ 19.42 $\downarrow$
Rephrase 60.32 $\uparrow$ 32.90 $\uparrow$ 78.50 $\uparrow$ 78.62 $\uparrow$ 18.90 $\downarrow$
Self-Output 60.05 $\uparrow$ 34.40 $\uparrow$ 80.12 $\uparrow$ 78.01 $\uparrow$ 19.88 $\downarrow$
$STM_{\tau=2.5}$ 61.64 $\uparrow$ 33.40 $\uparrow$ 75.09 $\downarrow$ 75.06 $\uparrow$ 20.27 $\downarrow$
4.1 In Domain and Out Domain Datasets

### v2

# Mitigating Forgetting in LLM Fine-Tuning via Low-Perplexity Token Learning

@@ Abstract
Maintaining consistent model performance across domains is a fundamental challenge in machine learning. While recent work has explored using LLM-generated data for fine-tuning, its impact on cross-domain generalization remains poorly understood. This paper presents a systematic analysis revealing that fine-tuning with LLM-generated data not only improves target task performance but also reduces non-target task degradation compared to fine-tuning with ground truth data. Through analyzing the data sequence in tasks of various domains, we demonstrate that this enhancement of non-target task robustness stems from the reduction of high perplexity tokens found in LLM-generated sequences. Following our findings, we showed that masking high perplexity tokens in ground truth training data achieves similar non-target task performance preservation, comparable to using LLM-generated data. Extensive experiments across different model families and scales, including Gemma 2 IT 2B, Llama 3 8B Instruct, and 3 additional models, agree with our findings. To the best of our knowledge, this is the first work to provide an empirical explanation based on token perplexity reduction to mitigate catastrophic forgetting in LLMs after fine-tuning, offering valuable insights for developing more robust fine-tuning strategies.
@@ 1 Introduction
Supervised fine-tuning (SFT) of large language models (LLMs) [26] has proven highly effective for enhancing their ability to follow novel instructions [37] and produce useful outputs across a wide range of tasks, including summarization and web querying [31, 22, 1, 16]. As LLMs are increasingly applied in specialized domains such as arithmetic and programming assistance [33, 23], practitioners often encounter substantial computational and data cleaning issues when adapting these models to gain improvements. These difficulties are exacerbated by performance saturation in smaller LLMs, where benchmark improvements on tasks like MATH [12] and MBPP [3] tend to plateau as model size decreases. Moreover, fine-tuning can risk degrading a model’s general capabilities [28, 20, 27, 4]. This raises a critical question: how can we efficiently fine-tune pre-trained, instruction-following LLMs for domain-specific tasks without compromising other skill sets such as arithmetic or common sense reasoning, especially when the original fine-tuning data is unavailable or computational resources are constrained?
Recent work has shown that fine-tuning with model-generated data can be highly effective, often outperforming training on ground truth data when using high-quality generated samples [42, 10, 29]. For example, Ren et al. [28] demonstrated that model-generated responses can surpass human annotations in downstream performance, particularly when using more capable models (e.g., GPT-4) to generate training data and mitigate degradation on general tasks. Although they attributed part of this success to the “familiarity” of generated responses, measured via low perplexity, it remains unclear why such data works well for instruction tuning. Specifically: (1) The regenerated response may differ contextually from the ground truth, and such contextual bias could affect the utility of low-perplexity training. (2) The “non-forgetting” impact of low-perplexity training on the performance of tasks not being fine-tuned, which maintains the practicality of the model, is still underexplored. (3) Since model size plays a role in distillation benefits, it is important to disentangle the effects of model scale and evaluate low-perplexity training using models of the same size.
Our work offers an empirical explanation, grounded in loss and perplexity reduction, for why LLM-generated data tends to cause less degradation in fine-tuning outcomes. Through systematic analysis, we first show that fine-tuning with different types of LLM-generated data (e.g., self-generated and rephrased responses) significantly mitigates degradation on non-target tasks compared to fine-tuning with ground truth data. Second, we trace this improvement to a key property illustrated in Figure 1: LLM-generated responses exhibit lower sentence-level perplexity and a smaller proportion of high-perplexity tokens. This leads to a practical insight: by simply masking high-perplexity tokens in ground truth training data, we can achieve a comparable reduction in non-target task degradation as observed with LLM-generated data. Importantly, this can be done using the same model prior to training, allowing us to pre-process the data and harness the benefits of low-perplexity training.
We propose a novel empirical explanation for non-target task performance degradation, introducing Selective Token Masking (STM), a simple yet effective strategy that filters out high-perplexity tokens to enable low-perplexity training. We have applied STM on (1) different perplexity criteria, (2) different model families and scales, and (3) various fine-tuning techniques from full-weight fine-tuning to parameter-efficient fine-tuning. We found that STM could achieve almost the same capabilities as the original models have on non-target tasks after fine-tuning on target tasks under all circumstances. We could identify that the degradation results from the low model weights update during the low-perplexity training. Together, these insights provide a new perspective on designing fine-tuning strategies that better retain general capabilities while enhancing performance on domain-specific tasks.
@@
Ground Truth: Denote the roots of $5x^{2}+3x+{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}4}$ by a and b. We have that $\alpha=\frac{1}{a}$ and $\beta=\frac{1}{b}$. So, $\alpha+\beta=\frac{1}{a}+\frac{1}{b}=\frac{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}a+}b}{ab}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}.}$ Now, we know that $a{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}+}b=\frac{-3}{5}$ and $ab=\frac{4}{5}$ by the relationship between sum/products of roots and the coefficients of a polynomial. Hence $\alpha+\beta=\dfrac{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}a}+b}{ab}={\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\boxed{-\dfrac{3}{4}}}$
Self-Output (Llama 3 8B Instruct): Let $r_{1}$ and $r_{2}$ be the roots of $45x^{2}+3x+$ 4. Then, by Vieta’s formulas, we have $r_{1}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}+r}_{2}=-\frac{3}{5}$ and $r_{1}r_{2}=\frac{4}{5}$. Let $\alpha=\frac{1}{r_{1}}$ and $\beta=\frac{1}{r_{2}}$. Then, we have $\alpha{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}+}\beta=\frac{1}{r_{1}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}+}\frac{1}{r_{2}}=\frac{r_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}2}}+r_{1}}{r_{1}r_{2}}=\frac{-\frac{3}{5}}{\frac{4}{5}}=\boxed{-\frac{3}{4}}$
Figure 1: An example MATH problem showing more high perplexity tokens (highlighted in red, perplexity $\geq 2.5$) in ground truth than Self-Output responses (Llama 3 self-generated responses).
@@ 2 Training LLMs on self-generated data
To evaluate the effectiveness of LLM-generated data for fine-tuning, we explore two distinct methods for generating training data across two different target tasks. We then assess the models fine-tuned with the generated data on five non-target tasks, respectively. The two generation strategies, Self-Output and Rephrase, offer complementary approaches to constructing LLM-based training data, each addressing different trade-offs and challenges. In this section, we introduce the original datasets used for data generation and describe our methodology for constructing self-generated training datasets using language models.
@@ 2.1 Training and Evaluation Framework
For target-task fine-tuning, we adopt the MBPP and MATH datasets. These datasets provide rich annotations, including full solutions, reasoning steps, and test cases, beyond just the final answers, allowing models to learn comprehensive task-solving behaviors. For evaluation, in addition to using the test sets from MBPP and MATH, we assess model performance on GSM8K [7], ARC-Challenge [6], and BIRD [18] to examine generalization to non-target tasks involving various forms of reasoning and generation. A detailed description of the datasets is provided in Section 4.
Table 1: Perplexity average and variance over answer sequences of MBPP and MATH training datasets on Llama 3-8B Instruct. The average sentence perplexity is calculated by averaging the sum of each sentence’s perplexity over the token perplexity score.
Data Method Avg. PPL
Ground Truth 4.83 (7.04)
MBPP Rephrase 1.69 (0.16)
Self-Output 1.16 (0.01)
Ground Truth 2.45 (0.81)
MATH Rephrase 2.11 (9.28)
Self-Output 1.34 (0.03)
@@ 2.2 Self-Output
Self-Output follows a similar high-level approach, similar to other synthetic data generation techniques [42, 28], where a high-quality synthetic dataset is created by sampling diverse responses from an LLM and applying a strict filtering process. For each training instance, we use a language model $M$ (e.g., Llama 3 8B Instruct) to generate $N$ (e.g., 32) distinct responses with a temperature setting of $T=0.7$. The generated responses are then filtered to retain only those that are semantically aligned with the ground truth, ensuring the quality and relevance of the synthetic data.
Self-Output excels at producing high-quality training data in domains with reliably verifiable outputs (e.g., programming tasks with objective ground truths). However, it remains constrained to settings where such verification is feasible and imposes significant computational overhead from the multiple generations and filtering required to identify valid outputs.
@@ 2.3 Rephrase
Yang et al. [41] proposed self-distillation. The process uses an instruction-finetuned LLM to rephrase ground-truth responses in its style. By providing both the instruction and ground truth to the target LLM, it generates semantically equivalent reformulations with ground truth responses. Compared to Self-Output, Rephrase is computationally efficient, as it requires only a single pass of generation without additional filtering. However, Rephrase is susceptible to hallucinations, as evidenced by the higher variance in token-level perplexity (see Table  1). Despite this limitation, Rephrase offers versatility, as it does not rely on verifiable ground truth and can be applied to a broader range of tasks.
Table 2 presents the performance comparison between Self-Output, Rephrase, and Baseline Fine-tuning on ground truth using different models in terms of model size and model series across both target task improvement and non-target task degradation rate. Both task improvement rate (TI) and degradation rate (BWT) are adopted from Backward Transfer rate from [14] and modified as follows:
$\mathbf{TI}=(a^{(train)}_{target}-a^{(original)}_{target})/a^{(original)}_{target}.$ (1)
$\mathbf{BWT}=\frac{1}{T-1}\sum_{i=1}^{T-1}(a^{(train)}_{i}-a^{(original)}_{i})/a^{(original)}_{i}.$ (2)
Given $T$ tasks composed of one target task and $T-1$ non-target tasks, TI is the target task performance improvement percentage after training on the target task. BWT refers to the average performance degradation percentage of the $T-1$ non-target task after training on the target task. The models to fine-tune in this paper refer to generic instruction following models before fine-tuning with target task data, e.g., Llama 3 8B Instruct [8], Gemma 2 IT 2B [32], Mistral 7B Instruct [15], and OLMo 2 7B series [24]. For target task performance, Self-Output demonstrates consistent positive TI in both MBPP and MATH. For the non-target task, the Self-Output method achieves better BWT scores than the baseline fine-tuning in most cases, indicating much less degradation on non-target task capabilities. For more details, such as performance comparison with other models or other minor findings for self-output and rephrase method, please refer to Appendix B.
Figure 2: Fine-tuned Llama 3 8B Instruct on MBPP labels generated by various LLMs. Backward Transfer (BWT) measures performance drop on non-target tasks, showing that Low-perplexity labels training generally leads to less degradation (upper left) than ground truth (bottom right).
@@
Figure 3: Comparison of token-level perplexity (PPL) distributions between human-annotated Ground Truth (top) and Llama 3 8B Instruct generations for Rephrase and Self-Output sequence (middle and bottom), where PPL of Self-Output data is low and with fewer spikes.
2.4 Analysis of perplexity across each sample
In this section, we analyze the perplexity of datasets generated by Self-Output and Rephrase to quantify their impact on model uncertainty and fine-tuning stability. We first collect the inference statistics of GPT-4o, GPT-4o mini, Gemma 2 IT 27B, Llama 3 8B Instruct, and Gemma 2 IT 2B from all three methods: Baseline Fine-tuning, Rephrase, and Self-Output strategy, along with their BWT, as Figure 3 shows that a higher perplexity of target training data could result in a higher degradation on non-target task. Second, we further collect the Llama 3 8B Instruct inference statistics from three kinds of training data: Baseline Finetuning (ground truth data), Rephrase, and Self-Output strategy. We evaluate the perplexity of responses by summing up the negative log probability of each token conditioned on its previous context in the sentence to measure the uncertainty of model outputs over the entire sequence. Specifically, the perplexity $PPL$ [^1] In this section, we use natural log instead of base 2 for computation convenience across a sequence is computed as:
$PPL=-\exp(\frac{1}{N}\sum_{i=1}^{N}P(w_{i}\mid w_{1},...,w_{i-1})),$ (3)
where $P(w_{i}\mid w_{1},...,w_{i-1})$ represents the conditional probability of token $w_{i}$ given its preceding context, and $N$ is the total number of tokens in the sequence.
Table 1 demonstrates that the Self-Output strategy achieves the lowest average perplexity of 1.16 in MBPP and 1.34 in MATH. Among the three approaches, Ground Truth exhibits the highest uncertainty with a mean perplexity of 4.83, and the perplexity of Rephrase is slightly higher than Self-Output. The lower perplexity of Self-Output suggests that LLMs are less likely to deviate far from their initial weights when fine-tuning on downstream tasks created by Self-Output. This characteristic maintains the model’s original performance on non-target tasks while effectively learning from the target task training data.
@@ 2.4 Analysis of perplexity across each sample
To investigate predictability at the token-level, we sample a single MATH response and visualize the token-level perplexity value. Figure 3 reveals distinct patterns across Ground Truth, Rephrase, and Self-Output responses. The first 20 tokens’ variance is 2.04 for Ground Truth, 2.31 for Rephrase, and 0.04 for Self-Output, which confirms our observation: Ground Truth exhibits frequent perplexity spikes of high amplitude (perplexity around 6-8), which reflects the natural variability in human problem-solving expressions. While Rephrase maintains similar peak amplitudes to Ground Truth, i.e., more regular spike patterns, its variability is between Ground Truth and Self-Outputs. Self-Output demonstrates remarkably consistent low token perplexity values rarely exceeding 2, indicating highly predictable token sequences. These patterns suggest that Self-Output generates more deterministic solutions, potentially beneficial for maintaining consistent problem-solving approaches, while Ground Truth sequences introduce much higher variations. To demonstrate the different content of tokens distribution between Ground Truth and Self-Output responses, please refer to Appendix C.
3 Low-perplexity token learning via selective token masking
@@
Our earlier analysis (Figure 3) reveals a notable discrepancy in token perplexity distributions between Self-Output generated data and other sources. This observation raises a central question: Is the superior performance of Self-Output methods primarily due to the lower perplexity of their generated tokens? To explore this hypothesis and design more efficient fine-tuning strategies, we propose Selective Token Masking (STM), a novel approach to supervised fine-tuning (SFT) that explicitly leverages token-level perplexity.
@@
The core idea behind STM is simple yet effective: we use an existing instruction-tuned model to compute token-level perplexities and mask ground truth tokens whose perplexity exceeds a predefined threshold $\tau$ during training. Our investigation yields two key insights: First, STM highlights the pivotal role of token perplexity in shaping fine-tuning performance. Second, it shows that the gains observed from Self-Output methods can be replicated through perplexity-based token filtering, suggesting that performance improvements stem more from low-perplexity token distributions than from the self-generation process itself. STM thus offers not only computational efficiency but also a principled framework for understanding model adaptation during fine-tuning.
@@
While our approach shares conceptual similarities with prior work such as Lin et al. [19], Ankner et al. [2], it differs significantly in both implementation and efficiency. Previous methods typically involved a two-stage process: training a reference model on high-quality data to identify learnable (high-perplexity) tokens, followed by pre-training a larger LLM on this curated subset. In contrast, STM achieves similar goals in a streamlined single-stage workflow by directly using an existing instruction-tuned model to compute token perplexities in a single forward pass. This design avoids the need to train or maintain any additional models, significantly reducing computational overhead while preserving the benefits of low-perplexity training. This makes STM a compelling choice for efficient fine-tuning rather than full-scale pretraining.
Importantly, STM is orthogonal to other training strategies. It can be applied alongside different data sources (e.g., ground truth or Self-Output) and training techniques (e.g., full-weight fine-tuning or LoRA) to alleviate non-target task performance degradation. In the following sections, we explore how STM complements these various settings and contributes to more robust fine-tuning outcomes.
4 Experimental Setting
To evaluate the effectiveness of our proposed method, we adopt benchmark datasets specifically selected to assess model robustness on both target and non-target tasks after fine-tuning. We include only datasets with verifiable ground truth, allowing us to filter out incorrect self-generated responses for a fair comparison with self-synthesized baselines and to ensure high data quality. Following recent findings [4] that LoRA [13] effectively mitigates performance degradation, we standardize our experiments by using LoRA for all fine-tuning. We then vary training strategies and data sources to study their impact on performance preservation. Details of our evaluation prompts and training configurations are provided in Appendix D. All implementation code and datasets are publicly available at: https://github.com/xxx/yyy.
Table 2: Performance comparison across different models and target tasks in terms of task improvement (TI) on target test set and backward transfer (the degradation percentage, BWT) of non-target test set, and the cost of data related process before training (e.g., generation of self-output, token PPL calculation and STM training criteria selection). Self-Output and our STM method have generally better BWT and TI, showing a better preservation of generalization capabilities after fine-tuning.
Model Target task Method BWT(%) TI(%) Cost (GPU hours)
Gemma 2 IT 2B MBPP Baseline Fine-tuning -38.19 -21.76 0
Self-Output -8.10   5.70 12 Hours
Rephrase -3.23 -4.69 30 Minutes
STMτ=2.5 (Ours)   0.42   0.00 5 Minutes
MATH Baseline Fine-tuning -36.68 -22.78 0
Self-Output -1.73   9.06 $\geq$ 2 Days
Rephrase -14.06 -28.83 39 Minutes
STMτ=2.5 (Ours) -2.93   7.83 8 Minutes
Llama 3 8B Instruct MBPP Baseline Fine-tuning -34.71 -2.23 0
Self-Output   3.09   1.55 16.8 Hours
Rephrase -5.32 -9.58 36.8 Minutes
STMτ=2.5 (Ours) -0.16   3.20 4.5 Minutes
MATH Baseline Fine-tuning -14.12 -17.83 0
Self-Output   0.31   9.55 $\geq$2 Days
Rephrase -1.09   4.78 29.3 Minutes
STMτ=2.5 (Ours) -0.30   6.37 7 Minutes
4.1 Target and non-target datasets